# SynPlay: Importing Real-world Diversity for a Synthetic Human Dataset

### v1

@@ Appendix B SynPlay Statistics
Figure 5: 456 virtual players in SynPlay created using Character Creator [47].
@@ C.2 Experimental setting
Image size applied in YOLOv8 training/inference. We use the image size of 1280$\times$1280 for all datasets except for COCO, which uses an image size of 640$\times$640. This decision simply takes into account the original image size of the datasets. Even after rescaling, the size range of human instances in the compared datasets remains similar. When using other models, i.e., Mask2Former in semantic segmentation tasks and retinaNet in data-scarce tasks, the image size/scaling recommended in the original settings was used.
Training Mask2Former without the large-scale jittering (LSJ) augmentation [19]. We did not use the default LSJ augmentation in training the Mask2Former segmentation models solely for performance reasons. In all cases, segmentation accuracy were found to be significantly lower when LSJ augmentation was used. LSJ augmentation, which greatly expands the range of image scaling, may not be suitable for aerial-view detection, which mainly includes small-sized human instances. This performance degradation with LSJ augmentation is also observed in [22], which is a reputable literature in the field of self-supervised learning.
Settings for PT-FT. When using PT-FT in the general tasks, training specifications, including training epochs and learning rate, did not differ between pre-training and fine-tuning. In data-scarce tasks, we follow all the settings of [43] as outlined in PTL, while leaving out the progressive component.
Settings for data-scarce tasks. For all experiments performed for data-scarce tasks including the scaling behavior study, we follow all the settings and experimental environments of [44].
@@ C.3 Quantitative measures
Fréchet Inception Distance (FID) [23]. We utilized the PyTorch implementation of FID in [42] with the default setup to assess the fidelity and diversity for all the training datasets involved in our experiments. We did not perform image scaling on the input for any dataset, and the final average pooling features were used to compute FID.
Proportion of nadir-view instances. An instance with an elevation angle greater than 71.57 relative to the UAV as is considered to be a nadir-view instance, representing the maximum elevation angle for Archangel* [44]. To identify nadir-view instances for Archangel, we utilized the dataset metadata, i.e., UAV position. Similarly, for Archangel*, we determined if an instance was an nadir-view instance based on the source instance, also using the dataset metadata. In the case of SynPlay, we computed the elevation angle for each instance using the absolute 3D coordinates of the instance and the UAV provided by SynPlay.
@@ D.3 SynPlay sample images
(1) red light, green light [Uncaptioned image]   [Uncaptioned image]   [Uncaptioned image] UGV UAV med-alt UAV low-alt [Uncaptioned image]   [Uncaptioned image]   [Uncaptioned image] CCTV back UAV high-alt CCTV side (2) sugar candy [Uncaptioned image]   [Uncaptioned image]   [Uncaptioned image] CCTV side UAV high-alt UGV [Uncaptioned image]   [Uncaptioned image]   [Uncaptioned image] UAV low-alt CCTV front CCTV back (3) tug-of-war [Uncaptioned image]   [Uncaptioned image]   [Uncaptioned image] UAV low-alt CCTV front UAV high-alt [Uncaptioned image]   [Uncaptioned image]   [Uncaptioned image] CCTV back UAV low-alt UGV
(4) marbles Refer to caption   Refer to caption   Refer to caption UGV CCTV front UAV med-alt Refer to caption   Refer to caption   Refer to caption CCTV side UAV med-alt UAV low-alt (5) stepping stones Refer to caption   Refer to caption   Refer to caption CCTV front UAV med-alt UAV high-alt Refer to caption   Refer to caption   Refer to caption UGV UAV med-alt CCTV side (6) squid Refer to caption   Refer to caption   Refer to caption UGV CCTV front UAV high-alt Refer to caption   Refer to caption   Refer to caption CCTV side UGV UAV low-alt

### v2

@@ Appendix B SynPlay Statistics
Figure 5: 456 virtual players in SynPlay created using Character Creator.
@@ C.2 Experimental setting
Image size applied in YOLOv8 training/inference. We use the image size of $\times1280$ for all datasets except for COCO, which uses an image size of $\times640$. This decision simply takes into account the original image size of the datasets. Even after rescaling, the size range of human instances in the compared datasets remains similar. When using other models, i.e., Mask2Former in semantic segmentation tasks and retinaNet in data-scarce tasks, the image size/scaling recommended in the original settings was used.
Training Mask2Former without the large-scale jittering (LSJ) augmentation [GGhiasiCVPR2021]. We did not use the default LSJ augmentation in training the Mask2Former segmentation models solely for performance reasons. In all cases, segmentation accuracy were found to be significantly lower when LSJ augmentation was used. LSJ augmentation, which greatly expands the range of image scaling, may not be suitable for aerial-view detection, which mainly includes small-sized human instances. This performance degradation with LSJ augmentation is also observed in [KHeCVPR2022], which is a reputable literature in the field of self-supervised learning.
Settings for PT-FT. When using PT-FT in the general tasks, training specifications, including training epochs and learning rate, did not differ between pre-training and fine-tuning. In data-scarce tasks, we follow all the settings of [YShenCVPR2023] as outlined in PTL, while leaving out the progressive component.
Settings for data-scarce tasks. For all experiments performed for data-scarce tasks including the scaling behavior study, we follow all the settings and experimental environments of [YShenArXiv2024].
@@ C.3 Quantitative measures
Fréchet Inception Distance (FID) [MHeuselNeurIPS2017]. We utilized the PyTorch implementation of FID in [Seitzer2020FID] with the default setup to assess the fidelity and diversity for all the training datasets involved in our experiments. We did not perform image scaling on the input for any dataset, and the final average pooling features were used to compute FID.
Proportion of nadir-view instances. An instance with an elevation angle greater than 71.57 relative to the UAV is considered to be a nadir-view instance, representing the maximum elevation angle for Archangel* [YShenArXiv2024]. To identify nadir-view instances for Archangel, we utilized the dataset metadata, i.e., UAV position. Similarly, for Archangel*, we determined if an instance was an nadir-view instance based on the source instance, also using the dataset metadata. In the case of SynPlay, we computed the elevation angle for each instance using the absolute 3D coordinates of the instance and the UAV provided by SynPlay.
@@ D.3 SynPlay sample images
(1) red light, green light [Uncaptioned image]   [Uncaptioned image]   [Uncaptioned image] UGV UAV med-alt UAV low-alt [Uncaptioned image]   [Uncaptioned image]   [Uncaptioned image] CCTV back UAV high-alt CCTV side (2) sugar candy [Uncaptioned image]   [Uncaptioned image]   [Uncaptioned image] CCTV side UAV high-alt UGV [Uncaptioned image]   [Uncaptioned image]   [Uncaptioned image] UAV low-alt CCTV front CCTV back
(3) tug-of-war [Uncaptioned image]   [Uncaptioned image]   [Uncaptioned image] UAV low-alt CCTV front UAV high-alt [Uncaptioned image]   [Uncaptioned image]   [Uncaptioned image] CCTV back UAV low-alt UGV (4) marbles [Uncaptioned image]   [Uncaptioned image]   [Uncaptioned image] UGV CCTV front UAV med-alt [Uncaptioned image]   [Uncaptioned image]   [Uncaptioned image] CCTV side UAV med-alt UAV low-alt
(5) stepping stones Refer to caption   Refer to caption   Refer to caption CCTV front UAV med-alt UAV high-alt Refer to caption   Refer to caption   Refer to caption UGV UAV med-alt CCTV side (6) squid Refer to caption   Refer to caption   Refer to caption UGV CCTV front UAV high-alt Refer to caption   Refer to caption   Refer to caption CCTV side UGV UAV low-alt
@@
Appendix E Benchmark Aerial View Dataset
Fig.13 presents qualitative detection results acquired on representative aerial-view human datasets, highlighting the fundamental differences between aerial and ground-view perception. Unlike conventional datasets such as MS COCO [TLinECCV2014], aerial human imagery involves small-scale human instances, often spanning only tens of pixels, and is captured from extreme and diverse viewpoints, including nadir, oblique, and off-nadir perspectives.
These factors create distinct challenges: while detailed appearance cues like facial features or textures become less informative at such scales, motion patterns, postures, and interaction dynamics remain critical. This requires datasets that concurrently capture both motion diversity and viewpoint variation.
SynPlay directly addresses these challenges by combining multi-perspective captures with rule-guided behavioral diversity, enabling robust human identification in aerial scenarios where existing datasets fall short as shown in the qualitative comparisons.
Refer to caption   Refer to caption   Refer to caption Refer to caption   Refer to caption   Refer to caption Refer to caption   Refer to caption   Refer to caption Refer to caption   Refer to caption   Refer to caption VisDrone [PZhuTPAMI2022] Okutama-action [MBarekatainCVPRW2017] SemanticDrone [ICGlink]
Figure 13: Qualitative detection samples for baselines vs. SynPlay+real models. Baseline models are trained on real dataset. $\square$ (green bounding boxes) indicate baseline (real only) human detection outputs while $\square$ (magenta bounding boxes) indicate the ones acquired by the models trained on corresponding SynPlay+real dataset. Note that unlike conventional ground-level benchmarks such as MS COCO [TLinECCV2014], aerial-view perception involves small-scale human instances, extreme viewpoint variations, and unique appearance challenges due to camera altitude and perspective. This highlights the need for specialized synthetic datasets like SynPlay that are tailored for long-range aerial human analysis.
Appendix F Limitations and Future Directions
A significant number of human instances in SynPlay appear at very low resolutions, which is an inherent challenge in any aerial-view dataset. However, as shown in the bounding box size histogram included in the supplementary material, SynPlay also contains many high-resolution human instances, with approximately 10,000 examples having bounding box areas greater than 10,000 pixels. We encourage selectively using these instances depending on the requirements of specific tasks.
SynPlay was developed to provide rich visual representations of human appearance for tasks focused on localizing people in complex scenes. While it centers on the human domain, we recognize the value of incorporating features from a broader set of object categories. Expanding SynPlay to include a wider array of objects could further enhance its utility for training and evaluating general-purpose models.
Future directions for the community include improving photorealism, simulating sensor artifacts such as rolling shutter and motion blur, and integrating synthetic and real data through hybrid training protocols. Adjusting scenario priors to support more diverse interactions and cultural settings, expanding metadata for occlusion and weather conditions, and incorporating additional sensing modalities like thermal imagery may also support greater robustness. More detailed behavioral annotations could further enable research in tracking, forecasting, and social interaction understanding. These extensions build on SynPlay’s strengths in controllability, scale, and behavioral diversity while helping shape the next generation of aerial perception benchmarks.